OpenStudy (anonymous):

police officer at rest at side of highway notices speeder moving at 17m/s along road.when speeder passes ,officer accelerates 3 m/ssquare in pursuit.speeder do not notice until police catch up. a)calculate speed of police car when officer catches him b)assume police car accelerates until its moving 10km/h faster than speeder and then moves at constant velocity until catching up with speeder.how long will it take to catch speeder?

OpenStudy (anonymous):

what's up @anitmary ?

OpenStudy (anonymous):

not so good with all this problems.........

OpenStudy (anonymous):

ok, so, did you get a)?

OpenStudy (anonymous):

ya its 33m/s......but didnt get b)??

OpenStudy (anonymous):

ok, looks like the easiest way is...

OpenStudy (anonymous):

the concept is \[d_{cop}=d_{speeder} ryt??\]

OpenStudy (anonymous):

yeah correct:)

OpenStudy (anonymous):

d=192.1 m & t=11.3 s

OpenStudy (anonymous):

for b) ?

OpenStudy (anonymous):

not for b).........m telling how i got a)....

OpenStudy (anonymous):

yep that's right for a)

OpenStudy (anonymous):

10km/h = 2.78 m/s so calculate the time and distance he covers getting up to speed time : Vf = Vi +at 19.78 = 0 +3t distance: 19.78^2 = 0^2 +2*3*d

OpenStudy (anonymous):

didnt you get time already?????????

OpenStudy (anonymous):

next find the distance the speeder covers in that time: (easymode) 17*t then see how far the officer has to go still in order to catch him... \[\huge 0 = d _{speeder} -d _{cop} -17t _{2} +19.78t _{2}\]

OpenStudy (anonymous):

sry, equation editor is bugging, might have to restart browser.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

then add the time you got from part one (the acceleration phase) to the time in part two (t2)

OpenStudy (anonymous):

t=6 s we got that ryt??

OpenStudy (anonymous):

i didnt understand your equation......

OpenStudy (anonymous):

which?

OpenStudy (anonymous):

0=dspeeder-dcop-17t2......??

OpenStudy (anonymous):

yeah I guess I wrote it weird try this... you got a distance for the officer and the speeder in the acceleration phase right?

OpenStudy (anonymous):

so how far is the speeder still ahead at that point?

OpenStudy (anonymous):

didnt understand wat happened after getting t=6s???

OpenStudy (anonymous):

so calculate the time and distance he covers getting up to speed time : Vf = Vi +at 19.78 = 0 +3t distance: Vf^2 = Vi^2 +2ad 19.78^2 = 0^2 +2*3*d

OpenStudy (anonymous):

t= 19.78/3 d=(19.78)^2 / 6

OpenStudy (anonymous):

meanwhile, the speeder has gone 17*(19.78/3) meters...

OpenStudy (anonymous):

so he's ahead by a distance of 17*(19.78/3) - (19.78)^2 /6

OpenStudy (anonymous):

agree so far?

OpenStudy (anonymous):

ohhkk....

OpenStudy (anonymous):

does it make sense? or is this just blowing your mind:)

OpenStudy (anonymous):

so even wen cop's initial speed is 10 km/h faster than speeder at t=6s ,he didnt catch up with him ryt??

OpenStudy (anonymous):

nope. not yet.. he's still behind a bit.

OpenStudy (anonymous):

yea..ok,got it..

OpenStudy (anonymous):

so, ready for the second bit? when he travels at a constant speed catching the rest of the way up to the speeder?

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

ok so 17*(19.78/3) - (19.78)^2 /6 + 17*t2 = 19.78*t2

OpenStudy (anonymous):

t2 is the time for the second part, the part with no acceleration.

OpenStudy (anonymous):

yes ...ok..

OpenStudy (anonymous):

find t2, add it to the first time (acceleration phase) and that's the total time to catch up to the speeder.

OpenStudy (anonymous):

so final time is 9.006s??? but text answer is 24s??

OpenStudy (anonymous):

46.88 = 2.78t

OpenStudy (anonymous):

\[d _{cop}=66.67m and d _{speeder}=17\times6.67=113.39m \]

OpenStudy (anonymous):

just solve this for t2: 17*(19.78/3) - (19.78)^2 /6 + 17*t2 = 19.78*t2 add it to the time in the first part.

OpenStudy (anonymous):

16.86 + 6.6

OpenStudy (anonymous):

ok i get 23....something.. i have a doubt........during that t=6s ,,,,vi=0 m/s & vf=20m/s ryt?? or was it vi=20 m/s and vf=0 m/s???

OpenStudy (anonymous):

Vf =19.78

OpenStudy (anonymous):

when there is almost 46.7 m to be covered(period of constant velocity) ,,vi=20 m/s and vf=0 m/s ryt??

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

there is no "Vf' or "Vi" for that part... both are just travelling at the same velocity throughout. driver 17 m/s officer 19.78 m/s

OpenStudy (anonymous):

ohh...yea...i forgot